OpenStudy (anonymous):

simplify f(x)= (4x^2+7x-2) / 4x-1. using complete sentences explain why f(1)=3, f(0)=2 and f(-1), yet f(1/4) is undefined. show work please:(

OpenStudy (anonymous):

\[f(\frac{1}{4})\] is undefined becase \[4(\frac{1}{4})-1=1-1=0\] so the denominator will be zero, and you cannot divide by 0

OpenStudy (anonymous):

\[f(1)=3\] because if you replace x by 1 you will get 3

OpenStudy (anonymous):

and in fact \[f(x)=\frac{4x^2+7x-2}{ 4x-1}=\frac{(x+2)(4x-1)}{4x-1}=x+2\] so this function basically says "add 2" but it is still undefined at x = 1/4 because the equality i wrote above is only true if \[x\neq \frac{1}{4}\]